OpenStudy (anonymous):

solve 6x-10y=12; -15x+25y=-30.

OpenStudy (texaschic101):

lets see if we can't make this easier....lets reduce the equations 6x - 10y = 12 becomes 3x - 5y = 6 -15x + 25y = -30 becomes -3x + 5y = -6 now if you were to add these... 3x - 5y = 6 -3x + 5y = -6 --------------add 0 + 0 = 0 There is an INFINITE MANY SOLUTIONS for this set of equations

OpenStudy (anonymous):

thank you this is what i have , but wasn't sure thanks again

OpenStudy (texaschic101):

no problem :) sometimes it is a good idea to check yourself, just to be on the safe side.